# Retrieval-Augmented Review Generation for Poisoning Recommender Systems

### v1

@@ Abstract
Recent studies have shown that recommender systems (RSs) are highly vulnerable to data poisoning attacks, where malicious actors inject fake user profiles, including a group of well-designed fake ratings, to manipulate recommendations. Due to security and privacy constraints in practice, attackers typically possess limited knowledge of the victim system and thus need to craft profiles that have transferability across black-box RSs. To maximize the attack impact, the profiles often remains imperceptible. However, generating such high-quality profiles with the restricted resources is challenging. Some works suggest incorporating fake textual reviews to strengthen the profiles; yet, the poor quality of the reviews largely undermines the attack effectiveness and imperceptibility under the practical setting.
To tackle the above challenges, in this paper, we propose to enhance the quality of the review text by harnessing in-context learning (ICL) capabilities of multimodal foundation models. To this end, we introduce a demonstration retrieval algorithm and a text style transfer strategy to augment the navie ICL. Specifically, we propose a novel practical attack framework named RAGAN to generate high-quality fake user profiles, which can gain insights into the robustness of RSs. The profiles are generated by a jailbreaker and collaboratively optimized on an instructional agent and a guardian to improve the attack transferability and imperceptibility. Comprehensive experiments on various real-world datasets demonstrate that RAGAN achieves the state-of-the-art poisoning attack performance.
Index Terms:
adversarial learning, recommender systems, poisoning attacks, shilling attacks, neural networks, deep learning
I Introduction
@@
Existing studies satisfy only a subset of the above outlined criteria, preventing them from forming a comprehensive and robust solution. Conventional poisoning methods [22, 7, 16] primarily focus on optimizing attack effectiveness for specific collaborative filtering (CF) models [23], but often neglect stealthiness, making them more detectable. To mitigate this, recent studies have leveraged generative adversarial networks (GANs) [24] to produce more inconspicuous perturbations. However, excessive emphasis on imperceptibility may compromise the transferability of attacks [10]. To address this trade-off, subsequent works [11, 6] have integrated transferability optimization into the GAN framework, thereby improving both objectives simultaneously. By producing profiles containing only fake numerical ratings, these attacks have semantic gaps with contemporary RSs enhanced by textual reviews [25]. To solve this, we propose to incorporate textual reviews, which is a publicly available natural feature and contains rich semantic information reflecting user behaviors and item characteristics [26], to reinforce attack profiles, as demonstrated in the previous work R-Trojan [21]. Nevertheless, the generated reviews are of low quality, characterized by logical gaps, incoherence, factual inaccuracies, sparse content, and so on, rendering them noticeable and weakening the effectiveness of transferability optimization. Moreover, R-Trojan makes an unrealistic assumption that the full data can be accessed.
TABLE I: Comparison of data poisoning attack approaches, where $\mathcal{|U|}$ and $\mathcal{|V|}$ denote the number of users and items, respectively, $l$ is the length of recommendation list and $p$ represents the maximum percentage of the data that the attack requires. Note that several methods (e.g., TrialAttack and RegUP) propose multiple attacks that require different levels of knowledge of the target RS, and we choose the attacks require the least knowledge for comparison due to they are closest to real-world settings.
@@ II Related Works
II-A Unimodal Attacks on Recommender Systems
@@
Some data poisoning attacks are then proposed to optimize for specific types of RSs, such as PGA [7] for the MF-based, Co-visitation Injection [8] for the AR-based and a black-box poisoning strategy designed for the graph-based [9]. These attacks typically outperform conventional shilling attacks on the RSs they are tailored for. As shown in Table I, [8] and [9] explore the transferability of poisoning attacks, with a focus on the effectiveness of the attacks among traditional black-box RSs. Traditional CF methods are gradually replaced by non-linear deep learning (DL) paradigms [23], due to their restricted abilities to capture complex structure of massive interaction data. The typical approaches are NCF [35] and LightGCN [36]. However, there is a gap between existing traditional algorithm-specific attacks [7, 8, 9] and the growing prevalence of DL-based paradigms. To fill this gap, Huang et al. [16] propose a data poisoning attack tailored for DL-based RSs. Nevertheless, they demonstrate the transferability of the attack under unrealistic assumptions, as shown in Table I.
@@
II-B Multimodal Attacks on Recommender Systems
We propose R-Trojan [21] to mitigate the most above problems, which is a tailored transformer-powered GAN for generating attack profiles containing not only numerical ratings but also textual reviews. However, as illustrated in Table I, R-Trojan makes an unrealistic assumption of full access to the learning data, and relies on the fine-tuned GPT-2 [38] with simple prompts (i.e., topics + sentiments) to generate fake textual reviews. However, GPT-2-generated text suffers from issues such as factual inaccuracy, repetitive phrasing, semantic confusion and logical inconsistencies, leading to the low-quality profiles. To solve the problems, RAGAN introduces a more efficient and effective generation scheme. By harnessing powerful FMs [39] that pre-trains on extensive and diverse data and is built upon flexible and advanced transformers, it makes the reviews semantically meaningful, coherent and natural, thereby improving the transferability and imperceptibility of attack profiles under a more practical setting.
III Problem Formulation
@@ III-A Preliminaries and Notations
III-B Threat Model
III-B1 Attacker’s Objective
As discussed in Section I, we consider two practical properties of an RS attack method.
@@ III-B1 Attacker’s Objective
    Transferability. Let $h(t)$ to denote the hit ratio of a target item $t$, which is the percentage of normal users whose the recommendation lists include $t$ after the attack, i.e., $h(t)=\sum_{u\in\widehat{\mathcal{U}}_{t}}\frac{\mathbb{I}(t\in\widehat{\mathcal{V}}_{u,:k})}{|\widehat{\mathcal{U}}_{t}|}$, where $\mathbb{I}(\cdot)$ is an indicator function, $\widehat{\mathcal{U}}_{t}=\{u\in\mathcal{U}:r_{u,t}=0\}$ is the set of normal users (i.e., target users who have not yet interacted with $t$) and $\widehat{\mathcal{V}}_{u,:k}$ is the top-$k$ recommendation list of the normal user $u$. Our first objective is to maximize $h(t)$ on the black-box RSs.
@@
    Imperceptibility. Our second attack objective is to make our attack as imperceptible as possible to maximize the attack impact and the number of affected users.
III-B2 Attacker’s Knowledge
We assume that the attacker is able to access to $p$ percent of the training data $\mathcal{M}$ of the victim RS, where $p=1$ denotes full knowledge of data and $0<p<1$ indicates partial knowledge of data. This is realistic considering the openness nature [5] of a RS. Although our attack makes use of item metadata to reinforce the attack, such information is often publicly available, for example, the images/descriptions/titles on Amazon can be collected by directly browsing or writing crawlers. In addition, since real-world RSs are usually complex and flexible (e.g., they may adopt ensemble schemes and may be updated frequently), it is difficult to gain accurate knowledge (e.g., model architectures, parameters and implementation details) of the victim RSs [6]. We therefore consider black-box settings for the victim RSs.
III-B3 Attacker’s Capability
@@
III-C Formulate Attacks As An Optimization Problem
@@ IV RAGAN
IV-A Jailbreaker
Jailbreaker aims to generate transferable and imperceptible fake user profiles. Given that the review text is increasingly incorporated to improve the RSs [37] and informative textual reviews play an important role in explaining the reasons for the numerical ratings convincingly [21] and influencing users’ decisions [40], the profiles are created with fake numerical ratings and enhanced with their corresponding fake textual reviews. To this end, we introduce a behavior pattern learning scheme to produce fake numerical ratings. For the generation of the fake textual reviews, we propose a demonstration retrieval algorithm and a text style transfer strategy to harness the in-context learning capabilities of the multimodal FMs.
IV-A1 Fake Numerical Rating Generation
@@
Personalized Template Sampling. Some works (e.g., [15]) create attack profiles from scratch (i.e., noise), resulting in low-quality profiles being produced. To address this, inspired by [6, 11], we use benign user profiles containing real rating patterns as templates. Unlike these previous methods that randomly sample from $\mathcal{M}$, we prioritize user profiles based on three standards: 1) preference is given to profiles that have interacted with items similar to the target item (e.g., within the same category), as they better reflect realistic behaviors given that benign users typically do not interact with items arbitrarily (e.g., guitarists tend not to purchase drum-related products [21]) and are more susceptible to influence due to the users share similar tastes; 2) profiles exhibiting a rich interaction history are favored to enhance personalization; and 3) profiles containing the target item are explicitly excluded to align with target normal users. As such, these standards collectively improve the effectiveness and imperceptibility of the attack. Let $\mathcal{P}$ to denote the profile templates,
@@
IV-A2 Fake Textual Review Generation
Multimodal FMs have shown remarkable capabilities in language and vision understanding, reasoning and generation, revolutionizing a wide range of fields [45]. In-context learning empowers FMs to generalize to unseen downstream tasks purely through prompt-based conditioning, eliminating the need for parameter updates. Compared to fine-tuning and re-training approaches, which may be impractical due to restricted access to the model parameters (e.g., closed source models like GPT-4 and Claude 3) or constraints on computational resources required for large models such as LLaMA 3.3, such an effective and efficient paradigm gains increasing popularity [28]. As such, we carefully craft prompts to effectively harness the ICL capabilities of the FMs to improve the quality of the corresponding fake textual reviews. Enhancing review quality can strengthen the optimization procedure of the attack effectiveness, resulting in more effective attack profiles.
@@ IV-A2 Fake Textual Review Generation
IV-B Instructional Agent
It is not easy to access the accurate knowledge of the victim RSs and their detailed recommendation results in real-world scenarios [21]. As such, we introduce an instructional agent to measure how effective the attack is and then optimize the Jailbreaker to improve the attack transferability, as inspired by [6, 21]. Existing works [6, 18, 3, 13] often use rating-only-based RSs to evaluate the effectiveness, which is unsuitable for our profiles containing both numerical ratings and textual reviews. We thus adopt the review-based RS [25] to build the agent, as shown in Fig. 1(d), where the semantically rich textual reviews are introduced to enrich user and item features for accurately predicting user preferences, so that the valuable and effective instructions can be provided to the Jailbreaker within the limited accessible data.
@@ IV-B Instructional Agent
where $\delta_{u}$ and $\delta_{v}$ are, respectively, the set of historical reviews from $u$ and on $v$, and $q_{u}$ and $q_{v}$ indicate the user ID embeddings and the item ID embeddings, separately. $f_{\text{agent}_{\Theta}}(\cdot)$ is generally constructed with feature modeling and preference prediction components [25]. For example, in the architecture of the previous work DeepCoNN++[21], a TextCNN [37] is first used to extract features from the user and the item reviews. The features are then combined with the embeddings. Afterwards, the latent representations of the user network and the item network are concatenated and mapped to a shared feature space. Finally, a MLP followed by a Sigmoid is adopted for effectively conducting implicit recommendations.
IV-C Guardian
[…]
IV-D Learning
As discussed in Section III-C, we obtain high-quality attack profiles by solving a bi-level optimization problem.
TABLE II: Statistics of Datasets
@@ IV-D Learning
TABLE III: HR@10 and NDCG@10 of different attacks against various victim RSs on real-world datasets. We use bold fonts and asterisk symbols to denote the best performance and second best performance methods, respectively.
@@
The Upper-level Problem. $\mathcal{L}_{\text{trans}}$ aims to maximize $h(t)$. If $t$ is in the recommendation lists of the normal users, it is not necessary to optimize much. But if not, it is to minimize the prediction rating gap between $t$ and the predicted items that are in the recommended lists, so that the target item can be promoted to as many normal users as possible. We design the loss as follows.
@@
$\mathcal{L}_{\text{imper}}$ is weighted by $\lambda$ and combined with $\mathcal{L}_{\text{trans}}$ to form the final loss for optimizing the parameters of the generation module, as detailed in Section III-C. Inspired by the works [6, 21], the optimization procedure of RAGAN is detailed in Algorithm 1, where Adam is used as an optimizer.
V Experiment
V-A Experiment Setup
V-A1 Datasets Selection
We use three real-world datasets from different scenarios [50, 46] to evaluate RAGAN, as detailed in Table II. For Yelp [^3] https://www.yelp.com/dataset, we randomly select a subset to avoid exceeding the hardware limit. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation of RAGAN. We adopt the leave-one-out method [35] to select the training set and the test set, with a split ratio of 9:1.
V-A2 Baseline Attack Methods
We compare RAGAN with typical and state-of-the-art poisoning attacks including Random, Bandwagon [4], PGA [7], DCGAN [10], AUSH [11], DLA [16], RecUP [15], Leg-UP [6], TrialAttack [18] and R-Trojan [21]. The hyper-parameters of the baselines are set as suggested in the original papers.
TABLE IV: HR@10 and NDCG@10 of R-Trojan and RAGAN with different attack sizes against various review-based RSs on three real-world datasets, where the attacks are performed under the black-box and white-box settings, respectively.
Dataset Attack DeepCoNN DeepCoNN++
@@ V-A2 Baseline Attack Methods
V-A3 Targeted Recommender Systems
We consider three rating-only-based victim RSs, as existing works do [6, 11, 21]: WRMF [51], NCF [35] and LightGCN [36]. Moreover, to evaluate the attack effectiveness on review-based RSs under both black-box and white-box settings, we use DeepCoNN [26] and our surrogate model as victim RSs. The hyper-parameters of the victim RSs are set as suggested in the original papers.
V-A4 Evaluation Metrics
[…]
V-A5 Configuration of RAGAN
To maintain a fair comparison, we set the attack size to 3% of the population as default, which can clearly show the performance differences [6], and the profile size equals the average number of ratings per user in the data set. For RAGAN, we set ll=6, the number of in-context examples to 3, training epochs to 20, batch size to 256, learning rate to 0.001, $\lambda$ to 0.5 to balance the two attack goals. To improve the attack efficiency while reducing the attack costs, we adopt ‘gemini-1.5-flash’ for $f_{\text{FM}_{\zeta}}(\cdot)$, ‘vit_large_patch14_224_clip_laion2b’ for CLIP-ViT and ‘all-MiniLM-L6-v2’ for SBERT, where the hyper-parameters are configured as the default values provided by Hugging Face [^4] https://github.com/huggingface/transformers. For other hyper-parameters and the data pre-processing scheme, we follow the previous work R-Trojan [21] to ensure the effectiveness of the evaluation.
V-B Attack Transferability
V-B1 Overall Transferability
Table III and Table IV illustrate the impressive transferability of RAGAN, underscoring the critical role of high-quality textual reviews in strengthening the effectiveness of the attack profiles across diverse victim RSs. From Table III, we can find that LightGCN is more robust than NCF and WRMF, possibly due to its unique graph convolutions. For the review-based RSs, as shown in Table IV, DeepCoNN++ is more vulnerable than DeepCoNN, due to its internal knowledge can be accurately obtained by RAGAN. Furthermore, the results on Yelp exhibit less variability, which we attribute to its larger scale and higher density, making it more resistant to the perturbations.
V-B2 RAGAN vs. Attack Baselines
As shown in Table III and Table IV, RAGAN achieves the state-of-the-art performance among all evaluated methods against various victim RSs on real-world datasets, with the highest HR@10 and NDCG@10. By refining the quality of textual reviews, RAGAN demonstrates improved attack transferability compared to R-Trojan. R-Trojan greatly outperforms other baselines, with its key distinction being the incorporation of textual reviews, demonstrating that even basic utilization of the reviews can significantly improve the attack performance.
@@ V-B1 Overall Transferability
V-C Attack Imperceptibility
V-C1 Overall Imperceptibility

### v2

@@ Abstract
Recent studies have shown that recommender systems (RSs) are highly vulnerable to data poisoning attacks, where malicious actors inject fake user profiles, including a group of well-designed fake ratings, to manipulate recommendations. Due to security and privacy constraints in practice, attackers typically possess limited knowledge of the victim system and thus need to craft profiles that have transferability across black-box RSs. To maximize the attack impact, the profiles often remains imperceptible. However, generating such high-quality profiles with the restricted resources is challenging. Some works suggest incorporating fake textual reviews to strengthen the profiles; yet, the poor quality of the reviews largely undermines the attack effectiveness and imperceptibility under the practical setting. To tackle the above challenges, in this paper, we propose to enhance the quality of the review text by harnessing in-context learning (ICL) capabilities of multimodal foundation models. To this end, we introduce a demonstration retrieval algorithm and a text style transfer strategy to augment the navie ICL. Specifically, we propose a novel practical attack framework named RAGAN to generate high-quality fake user profiles, which can gain insights into the robustness of RSs. The profiles are generated by a jailbreaker and collaboratively optimized on an instructional agent and a guardian to improve the attack transferability and imperceptibility. Comprehensive experiments on various real-world datasets demonstrate that RAGAN achieves the state-of-the-art poisoning attack performance.
{IEEEImpStatement}
Recommender systems play a vital role across e-commerce, online content, and social media platforms, benefiting both users and businesses through personalized suggestions and improved engagement. These advantages also create incentives for malicious actors to exploit them. Recent studies reveal that modern recommender systems are vulnerable to data poisoning attacks, leading to unfair competition and loss of user trust. However, existing attack methods often have limited practicality, overestimating system robustness under real-world constraints. To address this, we introduce a retrieval-augmented review generation framework that improves both the effectiveness and efficiency of such attacks. Experiments on real-world datasets show that RAGAN uncovers hidden practical vulnerabilities across multiple AI-based recommenders, outperforming state-of-the-art methods by up to 70% in attack performance while producing more natural review text. These findings offer actionable guidance for developing robust, transparent, and trustworthy recommender systems, supporting safer digital marketplaces and fostering public trust in AI technologies.
{IEEEkeywords}
Adversarial Learning, Deep Learning, Large Language Models, Poisoning Attacks, Recommender Systems, Retrieval-augmented Generation
1 Introduction
@@
Existing studies satisfy only a subset of the above outlined criteria, preventing them from forming a comprehensive and robust solution. Conventional poisoning methods [22, 7, 16] primarily focus on optimizing attack effectiveness for specific collaborative filtering (CF) models [23], but often neglect stealthiness, making them more detectable. To mitigate this, recent studies have leveraged generative adversarial networks (GANs) [24] to produce more inconspicuous perturbations. However, excessive emphasis on imperceptibility may compromise the transferability of attacks [10]. To address this trade-off, subsequent works [11, 6] have integrated transferability optimization into the GAN framework, thereby improving both objectives simultaneously. By producing profiles containing only fake numerical ratings, these attacks have semantic gaps with contemporary RSs enhanced by textual reviews [25]. To solve this, a prior work, R-Trojan [21], proposes to incorporate textual reviews, which is a publicly available natural feature and contains rich semantic information reflecting user behaviors and item characteristics [26], to reinforce attack profiles. Nevertheless, the reviews generated by R-Trojan are of low quality, characterized by logical gaps, incoherence, factual inaccuracies, sparse content, and so on, rendering them noticeable and weakening the effectiveness of transferability optimization. Moreover, R-Trojan makes an unrealistic assumption that the full data can be accessed.
Table 1: Comparison of data poisoning attack approaches, where $\mathcal{|U|}$ and $\mathcal{|V|}$ denote the number of users and items, respectively, $l$ is the length of recommendation list and $p$ represents the maximum percentage of the data that the attack requires. Note that several methods (e.g., TrialAttack and RegUP) propose multiple attacks that require different levels of knowledge of the target RS, and we choose the attacks require the least knowledge for comparison due to they are closest to real-world settings. Note that ✓ = available; $\times$ = not available.
@@ 2 Related Works
2.1 Unimodal Attacks on Recommender Systems
@@
Some data poisoning attacks are then proposed to optimize for specific types of RSs, such as PGA [7] for the MF-based, Co-visitation Injection [8] for the AR-based and a black-box poisoning strategy designed for the graph-based [9]. These attacks typically outperform conventional shilling attacks on the RSs they are tailored for. As shown in Table 1, [8] and [9] explore the transferability of poisoning attacks, with a focus on the effectiveness of the attacks among traditional black-box RSs. Traditional CF methods are gradually replaced by non-linear deep learning (DL) paradigms [23], due to their restricted abilities to capture complex structure of massive interaction data. The typical approaches are NCF [35] and LightGCN [36]. However, there is a gap between existing traditional algorithm-specific attacks [7, 8, 9] and the growing prevalence of DL-based paradigms. To fill this gap, Huang et al. [16] propose a data poisoning attack tailored for DL-based RSs. Nevertheless, they demonstrate the transferability of the attack under unrealistic assumptions, as shown in Table 1.
@@
2.2 Multimodal Attacks on Recommender Systems
R-Trojan [21] has been proposed to mitigate the most above problems, which is a tailored transformer-powered GAN for generating attack profiles containing not only numerical ratings but also textual reviews. However, as illustrated in Table 1, R-Trojan makes an unrealistic assumption of full access to the learning data, and relies on fine-tuning equipped with very simple prompts (i.e., topics + sentiments) to generate fake textual reviews. These design choices make R-Trojan computationally expensive and impractical when model parameters are inaccessible (for instance, with state-of-the-art closed models such as GPT-5). In addition, the generated text often suffers from issues such as factual inaccuracy, repetitive phrasing, semantic confusion and logical inconsistencies [38], leading to the low-quality profiles and thus limited transferability and imperceptibility. To solve the problems, RAGAN introduces a more efficient and effective generation scheme. By harnessing ICL capabilities of multimodal FMs [39] that pre-trains on extensive and diverse data and is built upon flexible and advanced transformers, it produces reviews that are more semantically meaningful, coherent and natural, thereby improving the transferability and imperceptibility of attack profiles under a more practical setting (i.e., beyond black-box RSs, the setting involves partial data). Crucially, ICL enables flexible and scalable adaptation without any additional training or parameter updates, in contrast to fine-tuning.
3 Problem Formulation
@@ 3.1 Preliminaries and Notations
3.2 Threat Model
3.2.1 Attacker’s Objective
As discussed in Section 1, we consider two practical properties of an RS attack method.
@@ 3.2.1 Attacker’s Objective
    Transferability. Let $h(t)$ to denote the hit ratio of a target item $t$, which is the percentage of normal users whose the recommendation lists include $t$ after the attack, i.e., $h(t)=\sum_{u\in\widehat{\mathcal{U}}_{t}}\frac{\mathbb{I}(t\in\widehat{\mathcal{V}}_{u,:k})}{|\widehat{\mathcal{U}}_{t}|}$, where $\mathbb{I}(\cdot)$ is an indicator function, $\widehat{\mathcal{U}}_{t}=\{u\in\mathcal{U}:r_{u,t}=0\}$ is the set of normal users (i.e., target users who have not yet interacted with $t$) and $\widehat{\mathcal{V}}_{u,:k}$ is the top-$k$ recommendation list of the normal user $u$. Our primary objective is to maximize $h(t)$ on the black-box RSs.
@@
    Imperceptibility. Our secondary attack objective is to make our attack as imperceptible as possible to maximize the attack impact and the number of affected users.
3.2.2 Attacker’s Knowledge
We assume that the attacker is able to access to a fraction $p$ of the training data $\mathcal{M}$ of the victim RS, where $p=1$ denotes full knowledge of data and $0<p<1$ indicates partial knowledge of data. This is realistic considering the openness nature [5] of a RS. Although our attack makes use of item metadata to reinforce the attack, such information is often publicly available, for example, the images/descriptions/titles on Amazon can be collected by directly browsing or writing crawlers. In addition, since real-world RSs are usually complex and flexible (e.g., they may adopt ensemble schemes and may be updated frequently), it is difficult to gain accurate knowledge (e.g., model architectures, parameters and implementation details) of the victim RSs [6]. We therefore consider black-box settings for the victim RSs.
3.2.3 Attacker’s Capability
@@
3.3 Formulate Attacks As An Optimization Problem
@@ 4 RAGAN
4.1 Jailbreaker
Jailbreaker is a novel attack generator that aims to produce transferable and imperceptible fake user profiles. Unlike previous generators that manipulate ratings or reviews in isolation, Jailbreaker jointly models users’ numerical and textual behaviors to ensure realism. Given that textual reviews are increasingly incorporated to improve the RSs [37], and that informative reviews play an important role in explaining the reasons for the numerical ratings convincingly [21] and influencing users’ decisions [40], each attack profile is created with synthetic numerical ratings and enhanced with corresponding fabricated reviews. To this end, we introduce a behavior pattern learning scheme to produce fake numerical ratings. For the generation of the high-quality fake textual reviews, we propose a demonstration retrieval algorithm and a text style transfer strategy to harness the in-context learning capabilities of the multimodal FMs.
4.1.1 Fake Numerical Rating Generation
@@
Personalized Template Sampling. Some works (e.g., [15]) create attack profiles from scratch (i.e., noise), resulting in low-quality profiles being produced. To address this, inspired by [6, 11], we use benign user profiles containing real rating patterns as templates. Unlike these previous methods that randomly sample from $\mathcal{M}$, we prioritize user profiles based on three standards: 1) preference is given to profiles that have interacted with items similar to the target item (e.g., within the same category), as they better reflect realistic behaviors given that benign users typically do not interact with items arbitrarily (e.g., guitarists tend not to purchase drum-related products [21]) and are more susceptible to influence due to the fact that users share similar tastes; 2) profiles exhibiting a rich interaction history are favored to enhance personalization; and 3) profiles containing the target item are explicitly excluded to align with target normal users. As such, these standards collectively improve the effectiveness and imperceptibility of the attack. Let $\mathcal{P}$ to denote the profile templates,
@@
4.1.2 Fake Textual Review Generation
Multimodal FMs have shown remarkable capabilities in language and vision understanding, reasoning and generation, revolutionizing a wide range of fields [45]. In-context learning empowers FMs to generalize to unseen downstream tasks purely through prompt-based conditioning, eliminating the need for parameter updates. Compared to fine-tuning and re-training approaches, which may be impractical due to restricted access to the model parameters (e.g., closed source models like GPT-4 and Claude 3) or constraints on computational resources required for large models such as LLaMA 3.3, such an effective and efficient paradigm gains increasing popularity [28]. As such, we carefully craft prompts to effectively harness the ICL capabilities of the FMs to improve the quality of the corresponding fake textual reviews, as shown in Fig. 1(c). Enhancing review quality can strengthen the optimization procedure of the attack effectiveness, resulting in more effective attack profiles.
@@ 4.1.2 Fake Textual Review Generation
$\mathcal{I}^{\prime}=f_{d}(\mathcal{I}^{*})_{:k},$ (10)
[…]
4.2 Instructional Agent
It is difficult to obtain accurate knowledge of the victim RSs, their detailed recommendation results, and periodic feedback in real-world scenarios [21]. Moreover, frequent queries to the target system can raise suspicion and often require a large attack budget. As such, we introduce an instructional agent to measure how effective the attack is and then optimize the Jailbreaker to improve the attack transferability and stealth while reducing the budget under black-box settings, as inspired by [6, 21]. Existing works [6, 18, 3, 13] often use rating-only-based RSs to evaluate the effectiveness, which is unsuitable for our profiles containing both numerical ratings and textual reviews. We thus adopt the review-based RS [25] to build the agent, as shown in Fig. 1(d), where the semantically rich textual reviews are introduced to enrich user and item features for accurately predicting user preferences, so that the valuable and effective instructions can be provided to Jailbreaker within the limited accessible data.
@@ 4.2 Instructional Agent
where $\delta_{u}$ and $\delta_{v}$ are, respectively, the set of historical reviews from $u$ and on $v$, and $q_{u}$ and $q_{v}$ indicate the user ID embeddings and the item ID embeddings, separately. $f_{\text{agent}_{\Theta}}(\cdot)$ is generally constructed with feature modeling and preference prediction components [25]. Following the surrogate architecture of the previous work [21], a TextCNN [37] is first used to extract features from the user and the item reviews. The features are then combined with the embeddings. Afterwards, the latent representations of the user network and the item network are concatenated and mapped to a shared feature space. Finally, a MLP followed by a Sigmoid is adopted for effectively conducting implicit recommendations.
Table 2: Statistics of Datasets
@@
Table 3: HR@10 and NDCG@10 of different attacks against various victim RSs on real-world datasets. We use bold fonts and asterisk symbols to denote the best performance and second best performance methods, respectively.
@@
4.3 Guardian
[…]
4.4 Learning
As discussed in Section 3.3, we obtain high-quality attack profiles by solving a bi-level optimization problem.
The Upper-level Problem. $\mathcal{L}_{\text{trans}}$ aims to maximize $h(t)$. If $t$ is in the recommendation lists of the normal users, it is not necessary to optimize much. But if not, it is to minimize the prediction rating gap between $t$ and the predicted items that are in the recommended lists, so that the target item can be promoted to as many normal users as possible. The loss is designed as follows.
@@ 4.4 Learning
$\mathcal{L}_{\text{imper}}$ is weighted by $\lambda$ and combined with $\mathcal{L}_{\text{trans}}$ to form the final loss for optimizing the parameters of the generation module, as detailed in Section 3.3. Inspired by the works [6, 21], the optimization procedure of RAGAN is detailed in Algorithm 1, where Adam is used as an optimizer.
Table 4: HR@10 and NDCG@10 of R-Trojan and RAGAN with different attack sizes against various review-based RSs on three real-world datasets, respectively.
Dataset Attack DeepCoNN Review-based Agent
@@
5 Experiment
5.1 Experiment Setup
5.1.1 Datasets Selection
We use three widely-used real-world datasets from different scenarios [50, 46] to evaluate RAGAN, which are Amazon Musical Instruments, Amazon Automotive, and Yelp, as detailed in Table 2. For Yelp [^3] https://www.yelp.com/dataset, we randomly select a subset to avoid exceeding the hardware limit. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation of RAGAN. We adopt the leave-one-out method [35] to select the training set and the test set, with a split ratio of 9:1.
5.1.2 Baseline Attack Methods
Following the latest work [21], we compare RAGAN with a series of representative poisoning attacks including conventional: Random and Bandwagon [4], algorithm-specific: PGA [7] and DLA [16], and GAN-based: DCGAN [10], AUSH [11], RecUP [15], Leg-UP [6], TrialAttack [18] and state-of-the-art R-Trojan [21]. To maintain a fair comparison, for all the attack methods, we set the attack size to 3% of the population as default, which can clearly show the performance differences [6], and the profile size equals the average number of ratings per user in the data set. Moreover, to further ensure fair and consistent evaluations, all attacks are optimized using our surrogate agent to avoid high query costs while maintaining stealthiness under practical black-box settings. For reproducibility, the hyper-parameters of the baselines are set as suggested in the original papers.
5.1.3 Targeted Recommender Systems
We consider three representative rating-only-based victim RSs, as existing works do [6, 11, 21]: WRMF [51], NCF [35] and LightGCN [36]. Moreover, to comprehensively evaluate the attack effectiveness on review-based RSs, we use typical DeepCoNN [26] and the state-of-the-art agent model [21] as victim RSs. For reproducibility, the hyper-parameters of the victim RSs are set as suggested in the original papers.
5.1.4 Evaluation Metrics
[…]
5.1.5 Configuration of RAGAN
We set ll=6, the number of in-context examples to 3, training epochs to 20, batch size to 256, learning rate to 0.001, $\lambda$ to 0.5 to balance the two attack goals. To improve the attack efficiency while reducing the attack costs, we adopt ‘gemini-1.5-flash’ for $f_{\text{FM}_{\zeta}}(\cdot)$, ‘vit_large_patch14_224_clip_laion2b’ for CLIP-ViT and ‘all-MiniLM-L6-v2’ for SBERT, where the hyper-parameters are configured as the default values provided by Hugging Face [^4] https://github.com/huggingface/transformers. For other hyper-parameters and the data pre-processing scheme, we follow the previous work R-Trojan [21] to ensure the effectiveness of the evaluation.
5.2 Attack Transferability
5.2.1 Overall Transferability
Table 3 and Table 4 illustrate the impressive transferability of RAGAN, underscoring the critical role of high-quality textual reviews in strengthening the effectiveness of the attack profiles across diverse victim RSs. From Table 3, we can find that LightGCN is more robust than NCF and WRMF, possibly due to its unique graph convolutions. For the review-based RSs, as shown in Table 4, the review-based agent is more vulnerable than DeepCoNN, mainly because RAGAN is optimized on an instructional agent with similar architectures and behavioral patterns. Furthermore, the results on Yelp exhibit less variability, which we attribute to its larger scale and higher density, making it more resistant to the perturbations.
5.2.2 RAGAN vs. Attack Baselines
As shown in Table 3 and Table 4, RAGAN achieves the state-of-the-art performance among all evaluated methods against various victim RSs on real-world datasets, with the highest HR@10 and NDCG@10. By refining the quality of textual reviews, RAGAN demonstrates improved attack transferability compared to R-Trojan. R-Trojan greatly outperforms other baselines, with its key distinction being the incorporation of textual reviews, demonstrating that even basic utilization of the reviews can significantly improve the attack performance.
(a) RAGAN on Amazon
(b) RAGAN on Yelp
Figure 6: The transferability of RAGAN with partial knowledge against various black-box RSs on real-world datasets.
5.3 Attack Imperceptibility
5.3.1 Overall Imperceptibility